# Revisiting Koehler’s experiment of measuring the ratio of the specific heats of air by self-sustained oscillations: a more concise theoretical interpretation

### v1

@@ Appendix A Details of our experimental setup
Figure A1 shows the details of our experimental setup. The key parameters are as follows: The volume of vessel A is $V_{0}=2.5\,\mathrm{L}$, The ball is a stainless steel ball with a diameter of 14.00 mm and a mass of 11.40 g. The inner diameter of the glass tube is 14.05 mm. The outer ends of the glass tube and the fine gas inlet tube are connected to vessel A using frosted glass surfaces to ensure airtightness while facilitating assembly and disassembly. The hole in the wall of the tube was drilled with a wide-angle drill, with an inner diameter of approximately 1.0 mm and an outer diameter of approximately 2.5 mm. As illustrated in the figure, the specific shape of the hole is not a necessary requirement but is merely intended to facilitate manufacturing. If adjustment of the hole size is required, a thin sheet with a smaller aperture can be used to block the original hole. The air pump we employed is an aquarium air pump, model SB-748, manufactured by Zhongshan SOBO Electric Appliance Co., Ltd. (China), with a power rating of 8 W and a maximum air output of 7 L/min. Vessel B serves as a gas reservoir, functioning to buffer and reduce pressure, thereby mitigating the effects of uneven air delivery from the air pump.
@@
Figure (A2) shows the variation of the air pressure inside the container over time during the self-sustained oscillations of the ball, measured by a miniature wireless Bluetooth barometer (SPL06-001 pressure sensor, Goertek Microelectronics Inc., China) placed inside the vessel A. The atmospheric pressure $B=92400\,\mathrm{Pa}$. It can be seen that the equilibrium pressure $P_{0}\sim 93155\,\mathrm{Pa}$, the period $T\sim 0.6\,\mathrm{s}$ and the angular frequency $\omega\sim 10\,\mathrm{rad/s}$. The dimensionless parameter $p_{0}=\frac{P_{0}}{B}=1.008$, $A=\frac{\pi RB}{m\omega^{2}}=1782$. The fluctuation of pressure within half a period $T/2$ is approximately $170\,\mathrm{Pa}$, so $\frac{\partial P}{\partial t}=jB-g(x)(P-B)$ is of the order of magnitude of $\frac{170\,\mathrm{Pa}}{0.3\,\mathrm{s}}$. Considering that the contributions from gas supply and leakage are comparable in scale, it is appropriate to assign gas leakage parameters
$e(P_{0}-B)\sim\frac{170\,\mathrm{Pa}}{0.3\,\mathrm{s}},$
$(e+f)(P_{0}-B)\sim 3\times\frac{170\,\mathrm{Pa}}{0.3\,\mathrm{s}}.$
Then we obtain
@@
the dimensionless
$G(y)=\begin{cases}G_{+}=(e+f)/\omega=0.25,\quad&y\geq 0,\\ G_{-}=e/\omega=0.075,\quad&y<0.\end{cases}$
As provided in our analysis section, for fixed point $(u^{*},p^{*})$,
@@
Let $J$ (gas supply parameter) as the control parameter, to generate self-sustained oscillations, the value of $J$ must be such that
$u^{*}_{+}=A\left[J-G_{+}(p_{0}-1)\right]<0$ (the ball’s motion tends towards the steady state of falling),
$u^{*}_{-}=A\left[J-G_{-}(p_{0}-1)\right]>0$ (the ball’s motion tends towards the steady state of rising).
Figure A2: The variation of the air pressure inside the container over time during the self-sustained oscillations of the ball, in an actual experiment.

### v2

@@ Acknowledgments
Author Contributions
All authors contributed equally to this work.
@@ Appendix A Details of our experimental setup
A key requirement for realizing the Koehler apparatus is achieving an appropriate match between the ball and the glass tube, with the difference in their diameters being only several tens of micrometers. Figure A1 shows the details of our experimental setup, which is based on a commercially available system HXWest . The ball is a bearing steel ball with a diameter of 14.00 mm and a mass of 11.40 g. The inner diameter of the glass tube is 14.05 mm. The tube and the fine gas inlet tube were connected to vessel A ($V_{0}=2.5\,\mathrm{L}$) through ground-glass joints, providing airtight sealing and convenient assembly. The hole in the wall of the tube was drilled with a wide-angle drill, with an inner diameter of approximately 1.0 mm and an outer diameter of approximately 2.5 mm. As illustrated in the figure, the specific shape of the hole is not a necessary requirement but is merely intended to facilitate manufacturing. If adjustment of the hole size is required, a thin sheet with a smaller aperture can be used to block the original hole. The air pump employed is an aquarium air pump (model SB-748, Zhongshan SOBO Electric Appliance Co., Ltd., China), with a rated power of 8 W, a maximum air flow rate of 7 L/min, and a price of less than \$10. Vessel B serves as a gas reservoir, functioning to buffer and reduce pressure, thereby mitigating the effects of uneven air delivery from the air pump.
@@
For constructing the experimental setup from individual components rather than using a pre-assembled commercial system, we recommend purchasing glass or acrylic tubes with a nominal inner diameter of 14 mm. Due to manufacturing tolerances, commercially available tubes often exhibit a slight positive deviation in inner diameter, which may be as large as approximately 14.10 mm in some cases. Two practical approaches can be used to achieve an appropriate ball–tube matching condition. The first approach is to directly select steel balls with a diameter close to the expected range (e.g., 14.15 mm), which can be used as received. The second approach is to start from a slightly oversized steel ball (e.g., 14.20 mm) and gradually reduce its diameter by polishing or fine grinding until a suitable match with the tube is obtained. The adequacy of the matching can be assessed experimentally by observing whether the ball can be lifted by the air flow within the tube.
@@
The air pressure inside the vessel was measured using a miniature wireless Bluetooth multisensor module (WT901BC, WitMotion Shenzhen Co., Ltd., China) WitMotion_WT901BC , costing approximately \$30. Besides air pressure, the module can also measure the three-axis magnetic field, acceleration, and angular velocity. This module incorporates an SPL06-001 barometric pressure sensor (Goertek Microelectronics Inc., China), with a relative accuracy of $\pm 0.06$ hPa and a typical absolute accuracy of $\pm 1$ hPa. The module was placed directly inside the vessel, as illustrated in Fig. A2.
Figure A2: Measurement of the pressure variation inside the container using a wireless pressure sensor. Compared with the setup in Fig. A1, a 2.5 L wide-mouth bottle was used as container A to accommodate the sensor. The container and glass tube were connected through an adapter plate, with the gaps sealed by hot-melt adhesive. The demonstration setup shown here was not finally sealed.
Figure A3 shows the temporal variation of the air pressure inside the container during the self-sustained oscillations of the ball. The atmospheric pressure was $P_{\mathrm{atm}}=92400~\mathrm{Pa}$. The equilibrium pressure was approximately $P_{0}\sim 93155~\mathrm{Pa}$, the oscillation period was $T\sim 0.6~\mathrm{s}$, and the corresponding angular frequency was $\omega\sim 10~\mathrm{rad/s}$. One can then determine the dimensionless parameters $p_{0}=P_{0}/P_{\mathrm{atm}}=1.008$, and $A=(\pi RP_{\mathrm{atm}})/(m\omega^{2})=1782$. The pressure variations within half a period is approximately $170\,\mathrm{Pa}$, so $\partial P/\partial t=jP_{\mathrm{atm}}-g(x)(P-P_{\mathrm{atm}})$ is of the order of magnitude of $567~\mathrm{Pa\,s^{-1}}$. Considering that the contributions from gas supply and leakage are comparable in magnitude, it is appropriate to assign gas leakage parameters
$e(P_{0}-P_{\mathrm{atm}})\sim 567~\mathrm{Pa\,s^{-1}},$
$(e+f)(P_{0}-P_{\mathrm{atm}})\sim 3\times 567~\mathrm{Pa\,s^{-1}},$
where the numerical factor $3$ is introduced as a representative choice for numerical estimation, without implying a specific relation between $e$ and $f$. Then we obtain
@@
and the dimensionless $G(y)$ is
$G(y)=\begin{cases}G_{+}=(e+f)/\omega=0.225,\quad&y\geq 0,\\ G_{-}=e/\omega=0.075,\quad&y<0.\end{cases}$
Then, from Eq. (13), the fixed point $(u^{*},p^{*})$ satisfies
@@
Let $J$ (the gas supply parameter) be taken as the control parameter. For self-sustained oscillations to occur, $J$ must satisfy
$u^{*}_{+}$ $=A\left[J-G_{+}(p_{0}-1)\right]<0,$
$u^{*}_{-}$ $=A\left[J-G_{-}(p_{0}-1)\right]>0.$
Figure A3: Measured time evolution of the air pressure inside the container during self-sustained oscillations of the ball. The mean value of the pressure corresponds to the equilibrium pressure $P_{0}$ and is shown with a horizontal line.